OpenStudy (anonymous):

Solve 4x2 = 12x + 40. factoring

OpenStudy (anonymous):

\[4x^2=12x-40\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I don't understand how you arrived at that answer....

OpenStudy (anonymous):

@jim_thompson5910 I don't understand how they got that.... Can you help me? I don't want to cheat....

OpenStudy (raden):

4x2 = 12x + 40 4x2 - 12x - 40 = 0 divide by 4 all sides, giving us x^2 - 3x - 10 = 0 factor out! (x-5)(x+2) = 0 setting each factor be equal 0 x-5 = 0 then x = 5 also x+2 = 0 then x = -2 so, the solution is {-2, 5}